# Overcoming Label Noise for Source-free Unsupervised Video Domain Adaptation

### v1

@@ IV-C Comparisons to the State-of-the-art Methods
To show that the efficacy of our proposed approach is not solely due to the addition of the motion stream with appearance, we show our adaptation results for both single-stream (appearance only) and two-stream (appearance and motion) models. Our single-stream model achieves 86.1% and 96.1% top-1 accuracy with a gain of 5.5% and 6.8% over the source-only model for UCF $\rightarrow$ HMDB and HMDB $\rightarrow$ UCF datasets respectively. Further improvement in the adaptation performance is observed when we couple our method CleanAdapt with the teacher-student framework [29, 30] resulting in CleanAdapt + TS. The teacher network serves as a regularization mechanism by producing consistent pseudo-labels, which, in turn, incentivizes the student model to make more confident predictions. This improved method CleanAdapt + TS achieves 88.6% and 96.7% top-1 accuracy, resulting in 8.0% and 7.4% gains over the source-only models, respectively.
In comparison, the best performing earlier existing model CoMix [16], which uses a temporal contrastive learning framework with background mixing, gives 6.4% and 5.1% gain for these two datasets, respectively. Note that all of these [11, 13, 16, 57, 32, 53, 55, 21, 14, 12, 15, 17, 33, 54, 56] methods use the source data along with the target data during adaptation, whereas we use only target data in our approach and attain similar gains.
@@
TABLE IV: Performance comparison with state-of-the-art video domain adaptation methods on EPIC-Kitchens dataset. Results with single-stream models are highlighted in cyan whereas the results with two-stream networks are highlighted in gray.
@@
MMD [7] ICML’15 I3D 43.1 48.3 46.6 55.2 39.2 48.5 46.8
AdaBN [82] PR’18 I3D 44.6 47.8 47.0 54.7 40.3 48.8 47.2
MCD [8] CVPR’18 I3D 42.1 47.9 46.5 52.7 43.5 51.0 47.3
MM-SADA [12] CVPR’20 I3D 48.2 50.9 49.5 56.1 44.1 52.7 50.3
STCDA [14] CVPR’21 I3D 49.0 52.6 52.0 55.6 45.5 52.5 51.2
Kim et al. [15] ICCV’21 I3D 49.5 51.5 50.3 56.3 46.3 52.0 51.0
MD-DMD [17] MM’22 I3D 50.3 51.0 56.0 54.7 47.3 52.4 52.0
CIA [54] CVPR’22 I3D 52.5 47.8 49.8 53.2 52.2 57.6 52.2
@@
DANN [6] ICML’15 I3D 38.3 38.8 37.7 42.1 36.6 41.9 39.2
ADDA [89] CVPR’17 I3D 36.3 36.1 35.4 41.4 34.9 40.8 37.4
TA3N [11] ICCV’19 I3D 40.9 39.9 34.2 44.2 37.4 42.8 39.9
CoMix [16] NeurIPS’21 I3D 38.6 42.3 42.9 49.2 40.9 45.2 43.2
@@
Liang et al. [22] ICML’20 Transformer 44.1 53.9 40.8 36.5 49.0 45.3 44.9
Li et al. [86] CVPR’20 Transformer 44.7 54.3 41.0 36.7 49.9 45.4 45.4
Kim et al. [51] TAI’21 Transformer 44.4 54.9 41.3 37.2 49.8 45.2 45.5
HCL [24] NeurIPS’21 Transformer 45.1 55.6 41.5 36.9 50.2 45.7 45.8
Huang et al. [33] MM’22 Transformer 46.3 58.2 42.2 38.1 52.3 46.1 47.2
@@
CleanAdapt + TS Ours I3D 48.3$\blacktriangle$ +6.5 48.7 $\blacktriangle$ +7.6 49.9 $\blacktriangle$ +8.0 56.3 $\blacktriangle$ +10.2 44.6 $\blacktriangle$ +7.3 48.9 $\blacktriangle$ +5.0 49.6 $\blacktriangle$ +7.6
@@
Similarly, our two-stream model CleanAdapt achieves state-of-the-art performance on both UCF $\rightarrow$ HMDB and HMDB $\rightarrow$ UCF datasets in terms of top-1 accuracy with the values of 89.8% and 99.2%, respectively. This is a significant gain of 7.3% for UCF $\rightarrow$ HMDB and 7.8% for HMDB $\rightarrow$ UCF over the source-only model without using any source-domain data which is much higher than the other source-dependent adaptation models [11, 13, 12, 14, 51, 16, 32, 17, 33, 21, 53, 54, 56, 55, 57]. Our improved method CleanAdapt + TS achieves a further gain of 11.1% for UCF $\rightarrow$ HMDB and 7.9%. This affirms the assertion that in source-free, unsupervised video domain adaptation, utilizing the low-loss samples from the target domain during the adaptation phase is justified. It also highlights the efficacy of employing a slowly updated teacher network for generating pseudo-labels to fine-tune the student network using strongly augmented target domain videos.
We now aim to show the effect of using high-loss samples for adaptation and discuss if overconfident pseudo-labels can affect the adaptation performance.
Do the overconfident pseudo-labels trigger error accumulation? A potential question arising here is whether the degradation of models and the further decline in the quality of pseudo-labels can occur due to the presence of overconfident yet incorrect pseudo-labels. Although error accumulation could be a possibility, we have found error accumulation to be negligible in practice. For example, the UCF pre-trained model selects low-loss samples with $\sim$98accuracy in each epoch of the adaptation stage from HMDB.
TABLE V: Performance comparisons with state-of-the-art video domain adaptation methods on UCF $\leftrightarrow$ HMDBsmall.
Method UCF normal-$\rightarrow$ HMDB HMDB normal-$\rightarrow$ UCF Source only 97.3 96.8 SHOT [22] 99.3 99.5 3C-GAN [86] 98.3 99.5 SFDA [51] 98.0 99.3 HCL [24] 99.3 99.5 MTRAN [33] 100 100 Source only 98.3 97.8 CleanAdapt + TS 100 100 Target Supervised 100 100
Comparisons with self-training based methods. In Table II, we compare our approach with the other self-training approaches [15, 16, 14]. Our method re-purposed the learning from noisy labels based pseudo-label selection method that shows better performance than all these.
@@
Figure 6: Class activation map (CAM) on target-domain videos of the UCF $\leftrightarrow$ HMDB dataset. The actions in green are correct predictions, while actions in red are incorrect. It is worth noting that the adapted model in the bottom row places greater emphasis on the action component rather than the contextual scene aspect. (Best viewed in color.)
@@
UCF normal-$\leftrightarrow$ HMDBsmall. To provide a thorough evaluation and compare fairly with existing methods, we compare the performance of our method CleanAdapt + TS on the UCF $\leftrightarrow$ HMDBsmall dataset in Table  V. The dataset size is very limited, and thus CleanAdapt + TS achieves scores comparable to the target-supervised baseline for both the UCF $\rightarrow$ HMDBsmall and HMDB $\leftrightarrow$ UCFsmall. We compare our approach CleanAdapt + TS with state-of-the-art techniques such as  [22, 51, 33, 86, 24] and achieves superior performance.
@@
EPIC-Kitchens. In Table IV, we compare the results of our approach with state-of-the-art image-based methods extended for videos [82, 7, 8, 6, 89, 22, 86, 24, 33] as well as video-based domain adaptation [12, 14, 15, 17, 54, 11, 16]. We implement our model to replicate the source-only and target-supervised performance as reported in [15]. Note that there is a minor difference ($\sim$2.7and $\sim$3.5%) in the performance of the source-only model reported in MM-SADA [12] and both of our models. Comparable distinctions to those observed in [12] can be identified in [16], primarily attributable to the non-deterministic operations associated with CUDA. However, such minor differences of source-only accuracy is not a concern for evaluating domain adaptation performance. The most important metric here is the gain achieved after adaptation over the source-only model. On average, the source-free methods [22, 86, 51, 24, 33] demonstrate a maximum improvement of 3% over the source-only model. In contrast, our improved method, CleanAdapt + TS, achieves 7.6% improvement despite being simple.
@@
MM-SADA [12] is the first to report domain adaptation results on the EPIC-Kitchens dataset, achieving an average of 4.8% gain on top of their source-only model followed by Song et al. [14] reporting an average gain of 5.7%. Kim et al. [15] show an improvement of 5.5% averaged over 6 datasets. However, all of these methods [12, 14, 15, 17, 54, 11, 16] use the source dataset for adaptation. In contrast to these prior approaches, our simple yet powerful source-free approaches, CleanAdapt and CleanAdapt + TS, achieve an average of 7.5% and 7.6% gain over the source-only model, respectively. The primary source of the performance boost achieved by our methods, despite their simplicity, can be attributed to the abundance of clean samples with low-loss values compared to the noisy ones.
@@
Figure 7: Hyperparameter search for the value of keep-rate $\tau$ for UCF101$\leftrightarrow$ HMDB51 and EPIC-Kitchens dataset. The keep-rate $\tau$ controls the number of samples to be selected as clean due to low-loss values computed against the pseudo-labels. The results reported here are for the two-stream network. (Best viewed in color.)
IV-D Hyperparameter Search
@@
TABLE VI: Cross-domain video retrieval results on UCF $\leftrightarrow$ HMDB dataset. Given queries from the target domain, we evaluate retrieved videos from the source domain in terms of R@k, where $k\in\{1,5,10\}$. Note that, all models reported here are two-stream networks and we average the similarity score from each modality to retrieve the source videos.
Method UCF normal-$\rightarrow$ HMDB HMDB normal-$\rightarrow$ UCF
@@
TABLE VII: Cross-domain video retrieval results on the EPIC-Kitchens dataset. Given queries from the target domain, we evaluate retrieved videos from the source domain in terms of R@k, where $k\in\{1,5,10\}$. All the models reported here are two-stream networks and we average the similarity score from each modality to retrieve the source videos.
Method D2 normal-$\rightarrow$ D1 D3 normal-$\rightarrow$ D1 D1 normal-$\rightarrow$ D2 D3 normal-$\rightarrow$ D2 D1 normal-$\rightarrow$ D3 D2 normal-$\rightarrow$ D3

### v2

@@ 4.3 Comparisons to the State-of-the-art Methods
To show that the efficacy of our proposed approach is not solely due to the addition of the motion stream with appearance, we show our adaptation results for both single-stream (appearance only) and two-stream (appearance and motion) models. Our single-stream model achieves 86.1% and 96.1% top-1 accuracy with a gain of 5.5% and 6.8% over the source-only model for UCF $\rightarrow$ HMDB and HMDB $\rightarrow$ UCF datasets respectively. Further improvement in the adaptation performance is observed when we couple our method CleanAdapt with the teacher-student framework [tarvainen2017mean, sohn2020fixmatch] resulting in CleanAdapt + TS. The teacher network serves as a regularization mechanism by producing consistent pseudo-labels, which, in turn, incentivizes the student model to make more confident predictions. This improved method CleanAdapt + TS achieves 88.6% and 96.7% top-1 accuracy, resulting in 8.0% and 7.4% gains over the source-only models, respectively.
[…]
Similarly, our two-stream model CleanAdapt achieves state-of-the-art performance on both UCF $\rightarrow$ HMDB and HMDB $\rightarrow$ UCF datasets in terms of top-1 accuracy with the values of 89.8% and 99.2%, respectively. This is a significant gain of 7.3% for UCF $\rightarrow$ HMDB and 7.8% for HMDB $\rightarrow$ UCF over the source-only model without using any source-domain data which is much higher than the other source-dependent adaptation models. Our improved method CleanAdapt + TS achieves a further gain of 11.1% for UCF $\rightarrow$ HMDB and 7.9%. This affirms the assertion that in source-free, unsupervised video domain adaptation, utilizing the low-loss samples from the target domain during the adaptation phase is justified. It also highlights the efficacy of employing a slowly updated teacher network for generating pseudo-labels to fine-tune the student network using strongly augmented target domain videos.
We now aim to show the effect of using high-loss samples for adaptation and discuss if overconfident pseudo-labels can affect adaptation performance.
Table 3: Comparison with state-of-the-art image-based source-free domain adaptation techniques. Method Backbone U $\rightarrow$ H H $\rightarrow$ U Source only TRN 72.7 72.2 Kim et al. [kim2021domain] TRN 69.9 74.9 Li et al. [li2020model] TRN 74.4 67.3 Yang et al. [yang2020unsupervised] TRN 75.3 76.3 Qiu et al. [qiu2021source] TRN 75.8 68.2 Source only I3D 80.6 89.3 Yang et al. [yang2021exploiting] I3D 86.6 91.4 Liang et al. [liang2020we] I3D 82.5 91.9 CleanAdapt I3D 86.1 96.1 CleanAdapt + TS I3D 88.6 96.7 Table 4: Performance comparisons with state-of-the-art video domain adaptation methods on UCF $\leftrightarrow$ HMDBsmall. Method U $\rightarrow$ H H $\rightarrow$ U Source only 97.3 96.8 SHOT [liang2020we] 99.3 99.5 3C-GAN [li2020model] 98.3 99.5 SFDA [kim2021domain] 98.0 99.3 HCL [huang2021model] 99.3 99.5 MTRAN [huang2022relative] 100 100 Source only 98.3 97.8 CleanAdapt + TS 100 100 Target Supervised 100 100
@@
Table 5: Performance comparison with state-of-the-art video domain adaptation methods on EPIC-Kitchens dataset. Results with single-stream models are highlighted in cyan whereas the results with two-stream networks are highlighted in gray. The average gains over the source-only model for the first and second best source-free approaches are highlighted in green and red, respectively.
@@
MMD [long2015learning] ICML’15 I3D 43.1 48.3 46.6 55.2 39.2 48.5 46.8
AdaBN [li2018adaptive] PR’18 I3D 44.6 47.8 47.0 54.7 40.3 48.8 47.2
MCD [saito2018maximum] CVPR’18 I3D 42.1 47.9 46.5 52.7 43.5 51.0 47.3
MM-SADA [munro2020multi] CVPR’20 I3D 48.2 50.9 49.5 56.1 44.1 52.7 50.3
STCDA [song2021spatio] CVPR’21 I3D 49.0 52.6 52.0 55.6 45.5 52.5 51.2
Kim et al[kim2021learning] ICCV’21 I3D 49.5 51.5 50.3 56.3 46.3 52.0 51.0
MD-DMD [yin2022mix] MM’22 I3D 50.3 51.0 56.0 54.7 47.3 52.4 52.0
CIA [yang2022interact] CVPR’22 I3D 52.5 47.8 49.8 53.2 52.2 57.6 52.2
@@
Source only I3D 35.5 38.1 39.4 40.5 32.0 39.2 37.5
CTAN [10139790] TCSVT’23 I3D 36.6 39.3 41.3 41.3 35.0 40.6 39.0
Target Supervised I3D 60.2 60.2 64.7 64.7 52.8 52.8 59.2
@@
DANN [ganin2015unsupervised] ICML’15 I3D 38.3 38.8 37.7 42.1 36.6 41.9 39.2
ADDA [tzeng2017adversarial] CVPR’17 I3D 36.3 36.1 35.4 41.4 34.9 40.8 37.4
TA3N [chen2019temporal] ICCV’19 I3D 40.9 39.9 34.2 44.2 37.4 42.8 39.9
CoMix [sahoo2021contrast] NeurIPS’21 I3D 38.6 42.3 42.9 49.2 40.9 45.2 43.2
@@
Liang et al[liang2020we] ICML’20 Transformer 44.1 53.9 40.8 36.5 49.0 45.3 44.9
Li et al[li2020model] CVPR’20 Transformer 44.7 54.3 41.0 36.7 49.9 45.4 45.4
Kim et al[kim2021domain] TAI’21 Transformer 44.4 54.9 41.3 37.2 49.8 45.2 45.5
HCL [huang2021model] NeurIPS’21 Transformer 45.1 55.6 41.5 36.9 50.2 45.7 45.8
MTRAN [huang2022relative] MM’22 Transformer 46.3 58.2 42.2 38.1 52.3 46.1 47.2 $\blacktriangle$ +3.0
@@
[…]
@@
Do the overconfident pseudo-labels trigger error accumulation? A potential question arising here is whether the degradation of models and the further decline in the quality of pseudo-labels can occur due to the presence of overconfident yet incorrect pseudo-labels. Although error accumulation could be a possibility, we have found error accumulation to be negligible in practice. For example, the UCF pre-trained model selects low-loss samples with $\sim$98% accuracy in each epoch of the adaptation stage from HMDB.
Comparisons with self-training based methods. In Table 2, we compare our approach with the other self-training approaches [kim2021learning, sahoo2021contrast, song2021spatio]. Our method re-purposed the learning from noisy labels based pseudo-label selection method that shows better performance than all these.
@@
UCF $\leftrightarrow$ HMDBsmall. To provide a thorough evaluation and compare fairly with existing methods, we compare the performance of our method CleanAdapt + TS on the UCF $\leftrightarrow$ HMDBsmall dataset in Table  LABEL:tab:uh-small. The dataset size is very limited, and thus CleanAdapt + TS achieves scores comparable to the target-supervised baseline for both the UCF $\rightarrow$ HMDBsmall and HMDB $\leftrightarrow$ UCFsmall. We compare our approach CleanAdapt + TS with state-of-the-art techniques such as  [liang2020we, kim2021domain, huang2022relative, li2020model, huang2021model] and achieves superior performance.
@@
EPIC-Kitchens. In Table 5, we compare the results of our approach with state-of-the-art image-based methods extended for videos as well as video-based domain adaptation. We implement our model to replicate the source-only and target-supervised performance as reported in [kim2021learning]. Note that there is a minor difference ($\sim$2.7% and $\sim$3.5%) in the performance of the source-only model reported in MM-SADA [munro2020multi] and both of our models. Comparable distinctions to those observed in [munro2020multi] can be identified in [sahoo2021contrast], primarily attributable to the non-deterministic operations associated with CUDA. However, such minor differences of source-only accuracy is not a concern for evaluating domain adaptation performance. The most important metric here is the gain achieved after adaptation over the source-only model. On average, the source-free methods demonstrate a maximum improvement of 3% over the source-only model. In contrast, our improved method, CleanAdapt + TS, achieves 7.6% improvement despite being simple.
@@
MM-SADA [munro2020multi] is the first to report domain adaptation results on the EPIC-Kitchens dataset, achieving an average of 4.8% gain on top of their source-only model followed by Song et al[song2021spatio] reporting an average gain of 5.7%. Kim et al[kim2021learning] show an improvement of 5.5% averaged over 6 datasets. CTAN [10139790] achieves a modest gain of 1.5% over the source-only models. However, all of these methods use the source dataset for adaptation. In contrast to these prior approaches, our simple yet powerful source-free approaches, CleanAdapt and CleanAdapt + TS, achieve an average of 7.5% and 7.6% gain over the source-only model, respectively. The primary source of the performance boost achieved by our methods, despite their simplicity, can be attributed to the abundance of clean samples with low-loss values compared to the noisy ones.
Table 6: Performance comparison of state-of-the-art video domain adaptation method on Ego2Exo [kalluri2024lagtran] dataset.
[…]
Unsupervised Adaptation
Source Only 8.39 15.66 12.03
TA3N [chen2019temporal] 6.92 27.95 17.44
TransVAE [wei2022unsupervised] 12.06 23.34 17.70
Zero-shot Video Recognition
EgoVLP [lin2022egocentric] 5.89 19.35 12.62
LaViLA [zhao2023learning] 5.86 19.13 12.50
LaGTran [kalluri2024lagtran] 12.34 30.76 21.55
Target Sup. 17.91 33.19 25.55
Source-free Unsupervised Adaptation
Source-only 18.08 40.02 29.05
CleanAdapt ($\tau=0.5$) 27.14 50.47 38.81
Target Sup. 30.99 52.05 41.52
Ego2Exo. Following [kalluri2024lagtran], we use the pre-extracted Omnivore Swin-L [girdhar2022omnivore] features in the Ego-Exo4D [grauman2024ego] dataset. We train a 2-layer linear classifier on top of this features keeping all other hyper-parameters as described above.
As shown in Table 6, TA3N [chen2019temporal] and TransVAE [wei2022unsupervised] achieve average improvements of 5.41% and 5.67%, respectively, over the source-only model on the Ego2Exo dataset. In contrast, the zero-shot baselines, EgoVLP [lin2022video] and LaViLA [zhao2023learning], deliver more modest gains of 0.59% and 0.47%, respectively. Although LaGTran [kalluri2024lagtran] achieves a 9.52% improvement over the source-only model, it relies on text data associated with the videos, which requires manual effort and contradicts our assumption. In contrast, our proposed method, CleanAdapt, achieves a 9.76% improvement over the source-only model without needing any text data in the target domain.
Figure 6: Class activation map (CAM) on target-domain videos of the UCF $\leftrightarrow$ HMDB dataset. The actions in green are correct predictions, while actions in red are incorrect. It is worth noting that the adapted model in the bottom row emphasizes the action component rather than the contextual scene aspect. (Best viewed in color.)
@@
Table 7: Performance comparison with zero-shot vision-language models on EPIC-Kitchens dataset. Results with zero-shot vision-language models are highlighted in cyan whereas the results our proposed model are highlighted in gray.
[…]
Zero-shot Baselines
Video-LLaVA [lin2023video] 32.64 32.64 33.73 33.73 34.90 34.90 33.76
EgoVLP [lin2022egocentric] 23.60 23.60 28.00 28.00 30.08 30.08 27.22
LaViLA [zhao2023learning] 46.44 46.44 50.13 50.13 47.64 47.64 48.07
Source-free Domain Adaptation
Source only 41.8 41.1 41.9 46.1 37.3 43.9 42.0
[…]
Target Supervised 62.3 62.3 72.7 72.7 71.1 71.1 68.4
Comparisons with Zero-Shot Vision-Language Models. In addition to the state-of-the-art video domain adaptation approaches, we also compare our methods CleanAdapt and CleanAdapt + TS with the following pre-trained vision-language models.
Video-LLaVA [lin2023video]. These baselines integrate visual representations into the language feature space, contributing to the development of unified Large Vision-Language Models (LVLMs). We prompt this LVLM with the following text - “USER: <video><video> Pick the action being performed in the video from the list below: [‘take’, ‘put’, ‘open’, ‘close’, ‘wash’, ‘cut’, ‘mix’, ‘pour’] ASSISTANT:”.
EgoVLP [lin2022egocentric]. A zero-shot baseline leverages the video-language pre-trained backbone from the Ego4D [grauman2022ego4d] dataset. We compute the embeddings for the class names using the distilbert-base-uncased model, and similarly, we extract video embeddings using the Ego4D [grauman2022ego4d] pre-trained backbone. Finally, the class name with the highest similarity is selected.
LaViLA [zhao2023learning]. This zero-shot baseline leverages the video-language pre-training by making a large-language model (LLM) conditioned on the visual inputs using the cross-attention layer on the Ego4D [grauman2022ego4d] dataset. Following [zhao2023learning], we compute the embeddings for the class names using the distilbert-base-uncased model, and similarly, we extract video embeddings using the Ego4D [grauman2022ego4d] pre-trained backbone. Finally, the class name with the highest similarity is selected.
The results of the zero-shot baselines on the EPIC-Kitchens dataset are presented in Table 7. From the table, it is evident that zero-shot vision-language models do not perform particularly well on egocentric video recognition tasks. For instance, Video-LLaVA [lin2023video], which is trained on generic third-person videos, achieves only 33.76% average top-1 accuracy. In contrast, EgoVLP [lin2022egocentric], and LaViLA [zhao2023learning], pre-trained on egocentric videos from the Ego4D [grauman2022ego4d] dataset, outperform Video-LLaVA but still fall short by 22.4% and 1.53% when compared with our proposed method CleanAdapt + TS.
@@
[…]
4.4 Hyperparameter Search
@@
4.5 Impact of Teacher-Student Framework
Table 8: Cross-domain video retrieval results on UCF $\leftrightarrow$ HMDB dataset. Given queries from the target domain, we evaluate retrieved videos from the source domain in terms of R@k, where $k\in\{1,5,10\}$. Note that, all models reported here are two-stream networks and we average the similarity score from each modality to retrieve the source videos.
Method UCF $\rightarrow$ HMDB HMDB $\rightarrow$ UCF
@@ 4.5 Impact of Teacher-Student Framework
4.6 Cross-domain Video Retrievals
In Table 8, we show the quantitative results for the cross-domain video retrieval task for the UCF $\leftrightarrow$ HMDB dataset. Our model retrieves better source videos from the target queries with R@1 of 0.92 and 0.91 as compared to the source-only model, which achieves only 0.82 and 0.88 on UCF $\rightarrow$ HMDB and HMDB $\rightarrow$ UCF datasets respectively. In Figure 9, we show qualitative retrieval results for the UCF $\rightarrow$ HMDB. Our model can correctly retrieve the source videos of the same semantic categories as the target query videos.
Table 9: Cross-domain video retrieval results on the EPIC-Kitchens dataset. Given queries from the target domain, we evaluate retrieved videos from the source domain in terms of R@k, where $k\in\{1,5,10\}$. All the models reported here are two-stream networks and we average the similarity score from each modality to retrieve the source videos.
Method D2 $\rightarrow$ D1 D3 $\rightarrow$ D1 D1 $\rightarrow$ D2 D3 $\rightarrow$ D2 D1 $\rightarrow$ D3 D2 $\rightarrow$ D3